# LinVideo: A Post-Training Framework towards $\mathcal{O}(n)$ Attention in Efficient Video Generation

### v2

@@ 2 Related Work
Efficient attention for video DMs. There have been lots of studies [7, 45] on video diffusion models (DMs) aiming to accelerate inference by sparsifying computationally expensive dense 3D attention. These methods fall into two categories: static and dynamic. Static methods [38, 60, 59, 22] predefine a sparsity pattern offline by identifying critical tokens. However, these methods fail to capture the dynamics of sparsity patterns during inference, leading to suboptimal performance. Dynamic methods [57, 48, 39, 61, 45] adapt the sparsity pattern at inference time based on input content. Thus, they need to select critical tokens through an additional identification step. Other approaches [63, 24] combine sparsification with quantization to further reduce latency and memory usage.
@@ 3 Preliminaries
In this work, we directly employ the effective kernel design from Hedgehog [58], which is formulated as:
@@ 4.2 Match the Distribution across Any Timestep
Table 1: Performance comparison with relevant baselines on 88 dimensions of VBench [17]. “+DMD2” denotes our 4-step distilled LinVideo model. We highlight the best score and the second score in bold and underlined formats, respectively.
@@
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Background Consistency $\uparrow$
Subject Consistency $\uparrow$
Scene Consistency $\uparrow$
Overall Consistency $\uparrow$
@@ 5.1 Experimental Details
Baselines. We compare LinVideo with sparse-attention baselines, including the static methods Sparse VideoGen (SVG) [45], Sparse VideoGen 2 (SVG2) [49], and DiTFastAttn (DFA) [55], and the dynamic method XAttention [48]. For latency, we include only the fast attention implementations of these methods to ensure a fair comparison, excluding auxiliary designs like the RMSNorm kernel in SVG. To be noted, we also exclude methods that require substantially more training resources [59] compared with LinVideo or support only specific video shapes [22]. Moreover, we do not compare our work with SLA [56], since its speedup relies on a specific CUDA kernel implementation that only supports the RTX 5090 GPU.
@@ 5.3 Ablation Studies
tar
Latency
$\downarrow$
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Overall Consistency $\uparrow$
@@
Method
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Overall Consistency $\uparrow$
@@
Method
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Overall Consistency $\uparrow$
@@
Figure 6: Training hours across different objectives. Settings are the same as those in Tab. 4.
Effect of ADM. We study several training objectives, as shown in Tab. 4. Our $\mathcal{L}_{\text{ADM}}$ outperforms the naive $\mathcal{L}_{\text{mse}}$ in Eq. (7) and the few-step distribution matching loss $\mathcal{L}_{\text{DMD}}$ that aligns only $p_{g}$ and $p_{0}$ (see Sec. 4.2), with a clear margin. In addition, $\mathcal{L}_{\text{ADM}}$ reduces training time by $\sim 4.4\times$ compared with $\mathcal{L}_{\text{DMD}}$ and the $\hat{s}_{t}^{\dagger}$ variant, both of which require training an extra model to estimate $\hat{s}_{t}$. Besides, our LinVideo transfers the model smoothly and progressively from quadratic- to linear-attention flow models. Thus, it is reasonable to view ${\bm{u}}_{\theta}$ as a flow model throughout this process. Experimentally, using ${\bm{u}}_{\theta}$ itself to compute $\hat{s}_{t}$ is more accurate, and thus achieves higher performance than $\hat{s}_{t}^{\dagger}$, which trains a separate score approximator.
@@ 6 Conclusion
  • Sec. H presents visualization examples of our LinVideo.
@@ Appendix C Ablation for $\alpha$
Range of $\alpha$
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Overall Consistency $\uparrow$
@@ Appendix F Directly combine Few-step Distillation with Linear Attention
Method
Imaging Quality $\uparrow$
Aesthetic Quality $\uparrow$
Motion Smoothness $\uparrow$
Dynamic Degree $\uparrow$
Overall Consistency $\uparrow$
@@ Appendix G Details of Heuristic Search
Appendix H Visualization Results

### v3

@@ 2 Related Work
Efficient attention for video DMs. There have been lots of studies [7, 50] on video diffusion models (DMs) aiming to accelerate inference [19] by sparsifying computationally expensive dense 3D attention. These methods fall into two categories: static and dynamic. Static methods [43, 65, 64, 26] predefine a sparsity pattern offline by identifying critical tokens. However, these methods fail to capture the dynamics of sparsity patterns during inference, leading to suboptimal performance. Dynamic methods [62, 53, 44, 66, 50, 31] adapt the sparsity pattern at inference time based on input content. Thus, they need to select critical tokens through an additional identification step. Other approaches [68, 28] combine sparsification with quantization [17, 18, 20] to further reduce latency and memory usage.
@@ 3 Preliminaries
In this work, we directly employ the effective kernel design from Hedgehog [63] [^1] Ablation study for kernel function can be found in Sec. I., which is formulated as:
@@ 4.2 Match the Distribution across Any Timestep
Table 1: Performance comparison with relevant baselines on 88 dimensions of VBench [21]. “+DMD2” denotes our 4-step distilled LinVideo model. We highlight the best score and the second score in bold and underlined formats, respectively. More results can be found in Sec. H and Sec. K.
@@ 5.1 Experimental Details
Baselines. We compare LinVideo with sparse-attention baselines, including the static methods Sparse VideoGen (SVG) [50], Sparse VideoGen 2 (SVG2) [54], and DiTFastAttn (DFA) [60], and the dynamic method XAttention [53]. For latency, we include only the fast attention implementations of these methods to ensure a fair comparison, excluding auxiliary designs like the RMSNorm kernel in SVG. To be noted, we also exclude methods that require substantially more training resources [64] compared with LinVideo or support only specific video shapes [26]. To be noted, in Sec. J, we compare our work with SLA [61], which relies on a specific CUDA kernel implementation that only supports the RTX 5090 GPU.
@@ 5.3 Ablation Studies
tar Latency (s) $\downarrow$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Overall Consistency $\uparrow$
@@
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Overall Consistency $\uparrow$
@@
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Overall Consistency $\uparrow$
@@
Figure 6: Training hours across different objectives. Settings are the same as those in Tab. 4. FLOPs comparison can be found in Tab. IV.
Effect of ADM. We study several training objectives, as shown in Tab. 4. Our $\mathcal{L}_{\text{ADM}}$ outperforms the naive $\mathcal{L}_{\text{mse}}$ in Eq. (7) and the few-step distribution matching loss $\mathcal{L}_{\text{DMD}}$ that aligns only $p_{g}$ and $p_{0}$ (see Sec. 4.2), with a clear margin. In addition, $\mathcal{L}_{\text{ADM}}$ reduces training time by $\sim 4.4\times$ compared with $\mathcal{L}_{\text{DMD}}$ and the $\hat{s}_{t}^{\dagger}$ variant, both of which require training an extra model to estimate $\hat{s}_{t}$. Besides, our LinVideo transfers the model smoothly and progressively from quadratic- to linear-attention flow models. Thus, it is reasonable to view $\hat{\bm{u}}_{\theta}$ as a flow model throughout this process. Experimentally, using $\hat{\bm{u}}_{\theta}$ itself to compute $\hat{s}_{t}$ is more accurate, and thus achieves higher performance than $\hat{s}_{t}^{\dagger}$, which trains a separate score approximator.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Natural Science Foundation of China (Nos. 62476018), and the Postdoctoral Fellowship Program of CPSF (No. BX20250487). This work was also supported by the Hong Kong Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R and NSFC/RGC Collaborative Research Scheme grant CRS_HKUST603/22.
@@ Acknowledgement
  • Sec. H provides additional comparison for DMD2 distilled few-step models;
  • Sec. I presents ablation studies for the kernel function of linear attention;
  • Sec. J provides a comparison between SLA and our method;
  • Sec. K shows the comparison results for CogVideoX [55];
  • Sec. L presents visualization examples of our LinVideo.
@@ Appendix C Ablation for $\alpha$
Range of $\alpha$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Overall Consistency $\uparrow$
@@ Appendix F Directly combine Few-step Distillation with Linear Attention
Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Overall Consistency $\uparrow$
@@ Appendix G Details of Heuristic Search
Appendix H Comparison for Few-step Models
Tab. IV  [^9] We report the average score of the chosen VBench dimensions Tab. 1. “FA2” denotes lossless FlashAttention2 baseline. shows that, when combined with DMD2, our method outperforms all baselines.
Table III: Comparison on Wan 1.3B (all methods use 4-step DMD2). Due to resource limits, we will add more results in the future.
Metric FA2 DFA XAttn SVG SVG2 LinVideo
[…]
Table IV: Training FLOPs ($3K$ steps).
[…]
LinVideo $5.97{\times}10^{4}$P
w/ $\mathcal{L}_{\text{mse}}$ $\mathbf{4.77{\times}10^{4}}$P
w/ $\mathcal{L}_{\text{DMD}}$ $2.97{\times}10^{5}$P
w/ $\hat{s}_{t}$ $2.97{\times}10^{5}$P
Table V: Comparison on CogVideoX-2B.
Metric FA2 DFA XAttn SVG SVG2 LinVideo
[…]
Table VI: Comparison with SLA. “*” denotes results on an RTX5090 GPU; other results are on an H100 GPU.
[…]
Wan 1.3B 67.63 97.32 162.65
SLA 65.72 93.95 78.42
LinVideo 67.61 68.26 112.14
Table VII: Comparison across kernel functions on Wan 1.3B.
[…]
Hedgehog 67.61 68.26
ReLU 65.48 68.12
Taylor Exp 67.24 68.54
Appendix I Ablation for Kernel Function
Hedgehog [63] preserves softmax-like spiky weights and dot-product monotonicity, making it more expressive than typical linear attention kernels. This benefits video generation (verify in Tab. VII  [^10] Taylor Exp: refer to Zhang et al. [63].), where long sequences require sharp, stable long-range interactions.
Appendix J Comparison between LinVideo and SLA
As shown in Tab. VII, LinVideo outperforms SLA on VBench. SLA’s large speedup mainly comes from its hardware-specific kernels (currently only supporting RTX5090), while LinVideo uses torch implementation. Moreover, LinVideo can be further accelerated if we combine sparse attention as SLA ($95\%$ sparsity).
Appendix K Comparison for CogVideoX
Our methods do not rely on any specific model architecture design. As shown in Tab. V, LinVideo outperforms baselines on CogVideoX.
Appendix L Visualization Results